Your oven is wired to 210V and draws 45A of current.

  1. What is the power of your oven?
  2. Power is usually sold using a price per the unit of kW-hr, which is 1000 Watts used for an hour. If you are baking a cake that takes 37 minutes to bake and the oven operates continuously, how much does it cost if Nova Scotia Power sells 1kW-hr for $0.156?
1
Expert's answer
2021-03-17T16:56:28-0400

1)

"P=VI=210\ V\cdot45\ A=9450\ W=9.45\ kW."

2)


"E=Pt=9.45\ kW\cdot37\ min\cdot\dfrac{1\ h}{60\ min}=5.83\ kWh,""Cost=5.83\ kWh\cdot 0.156\ \dfrac{\$}{kWh}=\$ 0.91."